OpenStudy (anonymous):

what is the proper equation in point-slope form for the given point and slope. Point: (5, 9); Slope: 2

OpenStudy (agreene):

Point slope form is defined as: \[y-y_1=m(x-x_1)\] So in this case. Assuming this is a (x,y) Cartesian pair... \[y-9=2(x-5)\]